A normal population has a mean of 11.2 and a standard deviation of 3.4. Refer to the table in Appendix B.1. a. Compute the z-value associated with 14.3. (Round the final answer to 2 decimal places.) z = b. What proportion of the population is between 11.2 and 14.3? (Round z-score computation to 2 decimal places and the final answer to 4 decimal places.) Proportion c. What proportion of the population is less than 10.0? (Round z-score computation to 2 decimal places and the final answer to 4 decimal places.) Proportion
Answer #1

Ans:

mean = 11.2

standard deviation =3.4

a)

z=(14.3-11.2)/3.4

z=0.91

b)

z(11.2)=0

z(14.3)=0.91

P(0<z<0.91)=P(z<0.91)-P(z<0)

=0.8186-0.5000=0.3186

c)

z=(10-11.2)/3.4

z=-0.35

P(z<-0.35)=0.3632